OpenStudy (dannyboi):

Find the 20th term of the following sequence. -6, -4, -2, 0, ... 1)32 2)44 3)46

OpenStudy (countryboyswag):

one way you can do this is count down from -6 by 2's til you get to the 20th number.

OpenStudy (eliesaab):

\[ a_1=-6\\ a_2=a_1+2=-4\\ a_3=a_2+2=a_1+2(2)=-2\\ \]

OpenStudy (dannyboi):

well whats the easiest way to do it?

OpenStudy (eliesaab):

What is \[a_n\]?

OpenStudy (dannyboi):

idk

OpenStudy (dannyboi):

well then

OpenStudy (eliesaab):

\[ a_n=a_1+(n-1)2=-6+2(n-1) \]

OpenStudy (eliesaab):

What is \[a_{20}\]?

OpenStudy (dannyboi):

i think i might be right but isnt it like 20a or something like that but its been a while since ive done algebra so i really dont know

OpenStudy (eliesaab):

\[a_{20} =-6 + 2(20-1)=-6+38=32 \]

OpenStudy (dannyboi):

ok